# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 4.4.2. Training
Full-parameter fine-tuning refers to tuning all parameters of a pre-trained language model for the downstream task of vulnerability detection. Tuning involves replacing the original output layer with a task-specific layer, such as a linear classifier layer, and training all initial model parameters (Liu et al., 2023b). This approach is widely adopted in the surveyed studies (Ahmad et al., 2021; Atiiq et al., 2024b, a; Bahaa et al., 2024; Buratti et al., 2020; Chan et al., 2023; Chen et al., 2024, 2023; Cipollone et al., 2025; Curto et al., 2024a, b; d’Aloisio et al., 2025; Deng et al., 2024; Ding et al., 2023, 2024b, 2025; Ehrenberg et al., 2024; Ferrag et al., 2025b; Fu et al., 2024; Fu and Tantithamthavorn, 2022; Grishina et al., 2023; Prabith et al., 2024; Gujar, 2024; Guo et al., 2024b; G”̈urfidan, 2024; Hanif and Maffeis, 2022; Huang et al., 2022; Ghafoor Hussain et al., 2025; Islam et al., 2024b; Jiang et al., 2024b, c; Kalouptsoglou et al., 2024a, b; Kim et al., 2022; Kong et al., 2024; Le et al., 2024; Li et al., 2021, 2025f; Lin et al., 2022a; Liu et al., 2024e, 2025c; Lu et al., 2023; Luo et al., 2021; Mamede et al., 2022; Manuel et al., 2024; Mechri et al., 2025; Mei et al., 2023; Ni et al., 2023b; Omar and Shiaeles, 2023; Pan et al., 2021; Purba et al., 2023; Qi et al., 2024; Rahman et al., 2024; Ren et al., 2024; Ridoy et al., 2024; Risse and B”̈ohme, 2024; Russo et al., 2025; Sahar et al., 2024; Shimmi et al., 2024b, a; Singh et al., 2022; Song et al., 2023; Steenhoek et al., 2025b, 2023b; Sultana et al., 2024; Sultan et al., 2024; Sun et al., 2024; Su and Wu, 2023; Thapa et al., 2022; Torkamani et al., 2025; Uddin et al., 2022; Wang et al., 2024b, c, 2023e; Wen et al., 2024a, b; Xiong and Dong, 2024; Xu et al., 2023; Yang et al., 2025a; Yin et al., 2024; Yuan et al., 2022; Yu et al., 2022; Zeng et al., 2023; Zhang et al., 2024d; Zhao et al., 2024a; Zheng et al., 2024; Zhou et al., 2024a, b; Zhu et al., 2022; Ziems and Wu, 2021; Akli et al., 2025; Halder et al., 2025; Ibanez-Lissen et al., 2025; Kaanan et al., 2024; Mächtle et al., 2025; Quan et al., 2025; Sun et al., 2023; Tian et al., 2025; Wen et al., 2023, 2025; Xue et al., 2025; Zheng et al., 2025; Lan et al., 2025; Panichella, 2025; Chen et al., 2025; Zhao and Guan, 2025). While this approach is computationally intensive and less reusable across tasks, it allows the model to specialize in vulnerability detection by adapting all internal representations.
@@
A specific variant of this technique is Instruction-Tuning, where a language model is fine-tuned on labeled input-output pairs annotated with natural language instructions (Du et al., 2024; Jiang et al., 2025; Yusuf and Jiang, 2024; Zhang et al., 2025b, 2024c; Bappy et al., 2025), e.g., ”Instruction: Detect whether the following code contains vulnerabilities” (Jiang et al., 2025). This method teaches the model to better follow human-written prompts by learning general patterns for task completion. In (Widyasari et al., 2025), Widyasari et al. use GPT-4o for generating the instructions. Instruction-tuning has shown promise in enhancing performance on unseen instructions and is often combined with parameter-efficient fine-tuning techniques to reduce training cost (Du et al., 2024; Jiang et al., 2025; Zhang et al., 2025b, 2024c).
Parameter-Efficient Fine-Tuning (PEFT)
@@
A key drawback of full-parameter fine-tuning is its high computational cost, which could potentially harm the model’s generalization ability. To address this drawback, several studies adopt parameter-efficient fine-tuning (PEFT) techniques, which adapt only a small subset of model parameters or learn external modules for new tasks while keeping most of the pre-trained model parameters frozen. To categorize the PEFT methods encountered in the surveyed studies, we adopt the taxonomy by Han et al. (Han et al., 2024), covering additive and reparameterized PEFT techniques.
Additive
@@
PEFT techniques introduce a small number of trainable parameters that are strategically positioned within the model architecture. During fine-tuning, only the weights of these additional modules or parameters are updated, which results in a substantial reduction in storage, memory, and computational resource requirements.
@@
Adapter-Tuning inserts small trainable neural modules, i.e., adapters, into a frozen pre-trained model. During fine-tuning, only the adapter parameters are updated, while the model’s parameters remain unchanged. Rather than employing default configurations that use the same adapter settings across all layers, Akli et al. (Akli et al., 2025) present AutoAdapt to automatically discover task-specific, layer-wide adapter configurations, allowing each layer to adopt distinct parameters.
@@
Prompt-Tuning prepends adjustable embedding vectors, referred to as soft prompts, to the start of the input sequence. This technique keeps the pre-trained model weights frozen while tuning only the parameters of the soft prompts (Chang et al., 2024; Lu et al., 2023; Wang et al., 2023e). Various adaptations of prompt-tuning have been investigated. Ren et al. (Ren et al., 2024) integrate prompt-tuning within a reinforcement learning framework. Feng et al. (Feng et al., 2025) explore graph-enhanced prompt-tuning, which incorporates structural code information by embedding graph features into soft prompt vectors.
@@
Apart from the methods mentioned above, Other-Additive approaches have been used that strategically train additional parameters during the fine-tuning process. Li et al. (Li et al., 2025b) introduce a vulnerability steering vector that represents the concept of vulnerability in the representation space. The vector is injected into the activation values of the corresponding layer, guiding the model’s behavior without modifying all its parameters. Wang et al. (Wang et al., 2024d) use (IA)3, an activation-scaling method that modifies intermediate activations using learnable rescaling vectors.
Reparameterized
@@
PEFT techniques transform a model’s architecture by modifying how its parameters are represented and trained. This typically involves replacing large components of the model with smaller, low-rank versions during training, reducing the number of parameters that need to be updated. At inference time, the model can be converted to its original parameterization.
@@
The most widely adopted reparameterization technique in the surveyed studies is categorized as Low-Rank Decomposition technique: LoRA (Low-Rank Adaptation) (Hu et al., 2022). LoRA injects trainable low-rank decomposition matrices into the attention layers of a frozen language model, enabling task-specific adaptation with significantly fewer parameters (Du et al., 2024; Jiang et al., 2025; Cao et al., 2024; Gonçalves et al., 2025; Guo et al., 2024b; Liu et al., 2025a; Luo et al., 2025; Mao et al., 2025; Shestov et al., 2025; Sultana et al., 2024; Wang et al., 2024e; Yang et al., 2025c; Zhou et al., 2024a; Zhu et al., 2024; Shu et al., 2025; Wen et al., 2025; Zhang et al., 2024c).
@@
LoRA Derivatives have also been explored to improve efficiency and adaptability. Tian et al. (Tian et al., 2024) use PiSSA (Principal Singular Values and Singular Vectors Adaptation) (Meng et al., 2024), which shares the same architecture as LoRA but adopts a different initialization method. Ibanez-Lissen et al. (Ibanez-Lissen et al., 2025) use GaLore (Gradient Low-Rank Projection) (Zhao et al., 2024c), a memory-efficient training strategy that allows full-parameter training while requiring less memory than common low-rank approaches such as LoRA. Another efficient PEFT design is QLoRA (Quantized LoRA) (Dettmers et al., 2023), which applies LoRA to a quantized version of the language model to further reduce memory consumption. QLoRA has been used in multiple surveyed studies to efficiently fine-tune large models for code vulnerability tasks (Zhang et al., 2025b; Curto et al., 2024a; Sultana et al., 2024; Yang et al., 2024a).
4.4.3. Learning Paradigms
@@
Beyond supervised learning, several advanced learning paradigms have been employed to enhance the performance and generalization capabilities of vulnerability detection models: contrastive learning, causal learning, multi-task learning, knowledge distillation, and continual learning.
Contrastive Learning
@@
Contrastive learning aims to learn more discriminative representations by encouraging the model to pull semantically similar instances closer and push dissimilar ones apart in the representation space, thereby improving the model’s ability to distinguish between structurally similar but semantically different code, such as secure versus vulnerable code. Several studies adapt contrastive learning techniques to the vulnerability detection domain (Ni et al., 2023b; Sahar et al., 2024; Wang et al., 2024e; Du et al., 2023; Ding et al., 2025; Ji et al., 2024). Notably, Du et al. (Du et al., 2023) propose mutual nearest neighbor contrastive learning to align the source and target domains. For multi-class classification, Ding et al. (Ding et al., 2025) introduce class-aware contrastive learning, which minimizes similarity only between samples with different class labels to improve type-specific discrimination. Ji et al. (Ji et al., 2024) present a hierarchical contrastive learning framework to bring vector representations of related CWEs closer together, combining supervised and self-supervised contrastive losses to promote geometric spread and better inter-class separation.
Causal Learning
@@
Causal learning guides models to focus on genuine cause-and-effect relationships rather than relying on spurious correlations in the data, e.g., variable or API names that are not actual causes of vulnerabilities. Rahman et al. (Rahman et al., 2024) identify such spurious features via semantic-preserving perturbations (e.g., variable renaming or dead-code injection) and then systematically remove their influence using causal inference techniques. Specifically, they simulate interventions, i.e., asking how the model would behave if certain spurious features were actively changed, and block non-causal paths by conditioning on known spurious factors. This strategy allows the model to rely on invariant, causally-relevant features and thereby improve robustness and generalization to unseen or perturbed code scenarios, such as different projects or naming conventions.
Multi-Task Learning
@@
Multi-task learning has emerged as an effective paradigm for improving generalization by jointly training on related tasks, thereby reducing overfitting and enhancing model robustness across a range of tasks. Several studies adopt multi-task learning to enhance vulnerability analysis. Fu et al. (Fu et al., 2024) fine-tune CodeBERT under the multi-task setting for predicting CWE-ID and CWE-type, while Chen et al. (Chen et al., 2025) combine CWE-ID classification with line-level vulnerability detection. Curto et al. (Curto et al., 2024b) perform binary vulnerability detection alongside multi-class vulnerability categorization. Similarly, Steenhoek et al. (Steenhoek et al., 2025b) fine-tune CodeBERT for simultaneous binary classification, multi-class vulnerability type prediction, and localization. Du et al. (Du et al., 2024) apply multi-task instruction-tuning for vulnerability detection, localization, and interpretation of root causes. Russo et al. (Russo et al., 2025) leverage shared information between self-admitted technical debt (SATD) and vulnerabilities to jointly detect both issues. Ding et al. (Ding et al., 2024b) use multi-task pre-training to jointly learn static and dynamic code properties.
Knowledge Distillation
@@
Knowledge distillation is a model compression strategy in which a smaller model, i.e., the student, is trained to replicate the behavior of a larger model, i.e., the teacher. By learning from the teacher’s output distributions, the student achieves comparable performance while significantly reducing computational cost, making it suitable for resource-constrained settings. Some studies directly adopt pre-trained distilled models such as DistilBERT for vulnerability detection (d’Aloisio et al., 2025; Liu et al., 2025a), while other studies apply knowledge distillation explicitly, using teacher and student models from the BERT family (Omar, 2023; Shi et al., 2024). Nguyen et al. (Nguyen et al., 2024) adopt a multi-teacher setup with CNN and GNN models to distill knowledge into a student LLM backbone. A more advanced approach is presented by Fu et al. (Fu et al., 2023a), which introduce a hierarchical knowledge distillation framework. The CWE label space is split into multiple sub-distributions based on semantic similarity, and individual CNN teacher models are trained on each subset. A student language model then learns to generalize across the teacher outputs, improving its ability to handle multi-class vulnerability detection tasks. Weyssow et al. (Weyssow et al., 2025) distill structured reasoning capabilities from a teacher LLM into smaller student LLMs using reinforcement learning from AI feedback (RLAIF). Panichella (Panichella, 2025) combines metamorphic testing with many-objective optimization for distillation of LLMs for code, measuring model robustness across semantically equivalent code variants. Ibanez-Lissen et al. (Ibanez-Lissen et al., 2025) use linear classifier probes (LPs) to identify valuable layers in pre-trained LLMs before fine-tuning, guiding selective compression and adaptation based on task-specific relevance.
Continual Learning
Retraining a model on all known vulnerabilities each time a new vulnerability is disclosed would demand considerable computational resources and time.
Continual learning enables models to incrementally learn new knowledge over time, such as new programming languages or up-to-date vulnerability data, without forgetting what has already been learned (Catastrophic Forgetting).
Gao et al. (Gao et al., 2023b) propose an adaptive regularization strategy that preserves the most representative (i.e., informative and diverse) samples in each dataset for model retraining under project-level, language-level, and time-level continual learning settings.
Similarly, Zhang et al. (Zhang et al., 2024d) present an incremental learning approach that allows the model to incorporate new programming languages without degrading performance on those it has already learned.
Tian et al. (Tian et al., 2025) introduce a training-free parameter fusion method that merges independent classification heads fine-tuned on different vulnerability types, allowing the fused model to adapt to new categories while preserving multi-class classification capabilities.
@@
Other data-centric learning techniques have been explored in the surveyed studies to improve model performance, i.e., Positive and Unlabeled (PU) learning, active learning, and adversarial training.
@@
Active learning aims to reduce labeling costs by strategically selecting the most informative samples for annotation. For example, Lan et al. (Lan et al., 2025) combine dataset maps with active learning to prioritize valuable samples while filtering out those that may harm performance.

### v2

@@ 4.4.2. Adaptation Technique (T2)
Full-Parameter Fine-Tuning (T2.2.2.1) refers to tuning all parameters of a pre-trained language model for the downstream task of vulnerability detection. Models are either fine-tuned by replacing the last output layer (Grishina et al., 2023) with a classification head (e.g., a multi-layer perceptron), which return the probability of vulnerability; or fine-tuned as a generative model where the response is matched to keywords such as “vulnerable” or “non vulnerable”. Full-parameter fine-tuning is widely adopted in the surveyed studies (Du et al., 2025b; Wu and Xiao, 2025; Liu et al., 2024a; Alam et al., 2025; Liu et al., 2026; Alrashedy et al., 2024; Ahmad et al., 2021; Aslan et al., 2025; Atiiq et al., 2024b, a; Bahaa et al., 2024; Buratti et al., 2020; Chan et al., 2023; Chen et al., 2024, 2023; Cheng et al., 2025; Cipollone et al., 2025; Curto et al., 2024b; d’Aloisio et al., 2025; Deng et al., 2024; Ding et al., 2023, 2024b, 2025; Ehrenberg et al., 2024; Ferrag et al., 2025b; Fu et al., 2024; Fu and Tantithamthavorn, 2022; Fu et al., 2023a; Gajjar et al., 2025a; Gao et al., 2023; Grishina et al., 2023; Prabith et al., 2024; Guo et al., 2024b; G”̈urfidan, 2024; Halder et al., 2025; Han and Zheng, 2025; Hanif and Maffeis, 2022; Huang et al., 2022; Ibanez-Lissen et al., 2025; Jiang et al., 2024c; JianJie and Le, 2023; Ji et al., 2024; Kaanan et al., 2024; Kalouptsoglou et al., 2024a, b; Kim et al., 2022; Kong et al., 2024; Lan et al., 2025; Le et al., 2024; Li et al., 2021, 2025k; Lin et al., 2022a; Li et al., 2025b, j; Liu et al., 2025a, 2024e, b; Lu et al., 2023; Luo et al., 2021; Mächtle et al., 2025; Maligazhdarova et al., 2025; Mamede et al., 2022; Mechri et al., 2025; Mei et al., 2023; Mock et al., 2025; Nguyen et al., 2025b; Ni et al., 2023b; Omar, 2023; Omar and Shiaeles, 2023; Panichella, 2025; Purba et al., 2023; Qi et al., 2024; Quan et al., 2025; Rahman et al., 2024; Ridoy et al., 2024; Risse and B”̈ohme, 2024; Russo et al., 2025; Safdar et al., 2025b; Shestov et al., 2025; Shi et al., 2024; Shimmi et al., 2024b, a; Simoni et al., 2025; Singh et al., 2022; Steenhoek et al., 2025b, 2023a, 2023b; Sultana et al., 2024; Sun et al., 2024; Su and Wu, 2023; Takieldeen et al., 2025; Thapa et al., 2022; Tian et al., 2025; Torkamani et al., 2025; Wang et al., 2025c, d, f; Wen et al., 2025, 2024a, 2024b, 2023; Xiong and Dong, 2024; Xu et al., 2023; Yang et al., 2025a; Yin et al., 2024; Yuan et al., 2022; Zeng et al., 2023; Zhang et al., 2024c; Zhao and Guan, 2025; Zhao et al., 2024a; Zheng et al., 2025, 2024; Zhou et al., 2024a; Zhu et al., 2025b, 2022; Ziems and Wu, 2021; Nguyen et al., 2025a), particularly with tiny models, see Figure 5. While full-parameter fine-tuning is computationally intensive and less reusable across tasks, it allows the model to specialize in vulnerability detection by adapting all internal representations.
@@
A specific variant of fine-tuning is Instruction-Tuning (T2.2.2.1.1), where a language model is fine-tuned on labeled input-output pairs annotated with natural language instructions (Bappy et al., 2025; Du et al., 2024; Islam et al., 2024b; Jiang et al., 2025b; Li et al., 2025c; Mao et al., 2025; Shu et al., 2025; Steenhoek et al., 2025a; Tian et al., 2024; Widyasari et al., 2025; Yusuf and Jiang, 2024; Zhang et al., 2025b, 2024b), e.g., ”Instruction: Detect whether the following code contains vulnerabilities” (Jiang et al., 2025b).
This method allows the model to better follow human-written prompts by learning general patterns for task completion.
In (Widyasari et al., 2025), Widyasari et al. use GPT-4o for generating the instructions.
Instruction-tuning has shown promise in enhancing performance on unseen instructions and is often combined with parameter-efficient (rather than full) fine-tuning techniques to reduce training cost (Du et al., 2024; Jiang et al., 2025b; Li et al., 2025c; Mao et al., 2025; Shu et al., 2025; Tian et al., 2024; Yusuf and Jiang, 2024; Zhang et al., 2025b, 2024b).
@@
Key drawbacks of full-parameter fine-tuning are its high computational cost and the potential harm to the model’s generalization ability. To address these drawbacks, several studies adopt Parameter-Efficient Fine-Tuning (PEFT) (T2.2.2.2) techniques, which adapt only a small subset of model parameters or learn external modules for new tasks while keeping most of the pre-trained parameters frozen. To categorize the PEFT methods, we adopt the taxonomy by Han et al. (Han et al., 2024), covering (i) selective, (ii) additive, and (iii) reparameterized PEFT techniques.
In Selective PEFT (T2.2.2.2.1), only a small subset of model parameters is selected for fine-tuning, while keeping the other parameters frozen. In the surveyed studies, the final (Zhang et al., 2025a) or upper model layer(s) (Chen et al., 2025b) were fine-tuned to generate contextual embeddings.
@@
Additive PEFT (T2.2.2.2.2) techniques introduce a small number of trainable parameters that are strategically positioned within the model architecture. During fine-tuning, only the weights of these additional modules or parameters are updated, reducing storage, memory, and computational resource requirements. Prominent additive PEFT techniques are adapter-tuning and prompt-tuning.
@@
Adapter-Tuning (T2.2.2.2.2.1) inserts small trainable neural modules, i.e., adapters, into a frozen pre-trained model. During fine-tuning, only the adapter parameters are updated, while the model’s parameters remain unchanged (Akli et al., 2025; Li et al., 2025e). Rather than employing default configurations that use the same adapter settings across all layers, Akli et al. (Akli et al., 2025) present AutoAdapt to automatically discover task-specific, layer-wide adapter configurations, allowing each layer to adopt distinct parameters.
@@
Prompt-Tuning (T2.2.2.2.2.2) involves learning prompts to improve model performance on a downstream task. Two main prompt forms are distinguished: hard prompts, which are manually crafted sequences of tokens appended to the input, and soft prompts, which are learned continuous embeddings that serve as virtual tokens prepended to the input. Prompt-tuning keeps the model weights frozen while tuning only the parameters of the soft prompts (Chang et al., 2024; Chen and Zhang, 2025; Feng et al., 2025; Lu et al., 2023; Ren et al., 2024). Various adaptations of prompt-tuning have been investigated. Ren et al. (Ren et al., 2024) integrate prompt-tuning with a reinforcement learning framework. Feng et al. (Feng et al., 2025) explore graph-enhanced prompt-tuning, which incorporates structural code information by embedding graph features into soft prompts.
@@
Apart from the methods mentioned above, another additive approach has been used that strategically trains additional parameters during the fine-tuning process (Additive-Other (T2.2.2.2.2.3)): Li et al. (Li et al., 2025d) introduce a vulnerability steering vector that represents the concept of vulnerability in the representation space. The vector is injected into the activation values of the corresponding layer, guiding the model’s behavior without modifying all its parameters.
@@
Reparameterized PEFT (T2.2.2.2.3) techniques transform a model’s architecture by modifying how its parameters are represented and trained. This typically involves replacing large components of the model with smaller, low-rank versions during training, reducing the number of parameters that need to be updated. At inference time, the model can be converted to its original parameterization (Han et al., 2024).
@@
The most widely adopted reparameterization technique in the surveyed studies is categorized as Low-Rank Decomposition (T2.2.2.2.3.1): LoRA (Low-Rank Adaptation) (Hu et al., 2022). LoRA injects trainable low-rank decomposition matrices into the attention layers of a frozen language model, enabling task-specific adaptation with significantly fewer parameters (Prasad et al., 2025; Alqarni and Azim, 2025; Aslan et al., 2025; Cao et al., 2024; Chen and Zhang, 2025; Curto et al., 2024a; Du et al., 2024; Gajjar et al., 2025b; Gonçalves et al., 2025; Guo et al., 2024b; Jararweh et al., 2025; Jiang et al., 2025b; Lekssays et al., 2025; Li et al., 2025j; Liu et al., 2025a; Li et al., 2025c; Luo et al., 2025; Mao et al., 2025; Shestov et al., 2025; Shu et al., 2025; Wang et al., 2024b; Wen et al., 2025; Weyssow et al., 2025; Yusuf and Jiang, 2024; Zhang et al., 2024b; Zhou et al., 2024a; Zhu et al., 2024).
@@
LoRA Derivatives (T2.2.2.2.3.2) have also been explored to improve efficiency and adaptability. Tian et al. (Tian et al., 2024) use PiSSA (Principal Singular Values and Singular Vectors Adaptation) (Meng et al., 2024), which shares the same architecture as LoRA but adopts a different initialization method. Ibanez-Lissen et al. (Ibanez-Lissen et al., 2025) use GaLore (Gradient Low-Rank Projection) (Zhao et al., 2024c), a memory-efficient training strategy that allows full-parameter training while requiring less memory than common low-rank approaches. Another efficient PEFT design is QLoRA (Quantized LoRA) (Dettmers et al., 2023), which applies LoRA to a quantized version of the language model to further reduce memory consumption. QLoRA has been used in multiple surveyed studies to efficiently fine-tune large models for code vulnerability tasks (Zhang et al., 2025b; Sultana et al., 2024; Yang et al., 2024a; Ouchebara and Dupont, 2025; Sun et al., 2025a; Wei et al., 2025).
Learning Paradigms (T2.3)
@@
Beyond standard supervised fine-tuning, several advanced learning paradigms have been employed to enhance the performance and generalization capabilities of vulnerability detection models, e.g., contrastive learning, causal learning, multi-task learning, knowledge distillation, continual learning, and reinforcement learning.
@@
Contrastive Learning (T2.3.1) aims to learn more discriminative representations by encouraging the model to pull semantically similar instances closer and push dissimilar ones apart in the representation space, thereby improving the model’s ability to distinguish between structurally similar but semantically different code, i.e., secure versus vulnerable code. Several studies adapt contrastive learning techniques to the vulnerability detection domain (Jiang et al., 2022; Wu and Xiao, 2025; Ni et al., 2023b; Wang et al., 2024b; Du et al., 2023; Ding et al., 2025; Ji et al., 2024; Gajjar et al., 2025a; Li et al., 2025e; Wang et al., 2025f). Notably, Du et al. (Du et al., 2023) propose mutual nearest neighbor contrastive learning to align the source and target domains. Ding et al. (Ding et al., 2025) introduce class-aware contrastive learning, which minimizes similarity only between samples with different class labels to improve type-specific discrimination. Ji et al. (Ji et al., 2024) present a hierarchical contrastive learning framework to bring vector representations of related CWEs closer together, combining supervised and self-supervised contrastive losses to promote geometric spread and improved inter-class separation.
@@
Causal Learning (T2.3.2) guides models to focus on genuine cause-and-effect relationships rather than relying on spurious correlations in the code, e.g., variable or API names that are not actual causes of vulnerabilities. Rahman et al. (Rahman et al., 2024) identify such spurious features via semantic-preserving perturbations (e.g., variable renaming or dead-code injection) and then systematically remove their influence using causal inference techniques. Specifically, they simulate interventions, i.e., asking how the model would behave if certain spurious features were actively changed, and block non-causal paths by conditioning on known spurious features. This strategy allows the model to rely on causally relevant features and, thereby, improve robustness and generalization to unseen or perturbed code scenarios, such as different projects or naming conventions.
@@
Multi-Task Learning (T2.3.3) has emerged as an effective paradigm for improving generalization by jointly training on related tasks, thereby reducing overfitting and enhancing model robustness across a range of tasks. Several studies adopt multi-task learning to enhance vulnerability analysis: Fu et al. (Fu et al., 2024) fine-tune under the multi-task setting for predicting CWE-ID and CWE-type, while Chen et al. (Chen et al., 2025b) combine CWE-ID classification with line-level vulnerability detection. Curto et al. (Curto et al., 2024b) perform binary vulnerability detection alongside multi-class vulnerability categorization. Similarly, Steenhoek et al. (Steenhoek et al., 2025b) fine-tune for simultaneous binary classification, multi-class vulnerability type prediction, and localization. Islam et al. (Islam et al., 2023) combine vulnerability classification with description. Du et al. (Du et al., 2024) apply multi-task instruction-tuning for vulnerability detection, localization, and interpretation of root causes; while Yang et al. (Yang et al., 2024a) focus on multi-task fine-tuning for vulnerability detection, explanation, and repair. Russo et al. (Russo et al., 2025) leverage shared information between self-admitted technical debt (SATD) and vulnerabilities to jointly detect both issues, Alqarni and Azim (Alqarni and Azim, 2025) combine the analysis of code, system calls, and bitstreams, and Ding et al. (Ding et al., 2024b) use multi-task pre-training to jointly learn static and dynamic code properties.
@@
Knowledge Distillation (T2.3.4) is a model compression strategy in which a smaller model, i.e., the student, is trained to replicate the behavior of a larger model, i.e., the teacher. By learning from the teacher’s output distributions, the student achieves comparable performance while significantly reducing computational cost, making it suitable for resource-constrained settings. Some studies directly adopt pre-trained distilled models such as DistilBERT for vulnerability detection (d’Aloisio et al., 2025; Liu et al., 2025a; Lin et al., 2022a), while other studies apply knowledge distillation explicitly, using teacher and student LLMs (Omar, 2023; Shi et al., 2024; Wei et al., 2025; Weyssow et al., 2025; Panichella, 2025; Liu et al., 2026), or a setup with CNN and GNN models to distill knowledge into or from a LLM backbone (Nguyen et al., 2025b; Liu et al., 2025b; Fu et al., 2023a; Han and Zheng, 2025). A more advanced approach is presented by Fu et al. (Fu et al., 2023a), who introduce a hierarchical knowledge distillation framework. The CWE label space is split into multiple sub-distributions based on semantic similarity, and individual CNN teacher models are trained on each subset. A student language model then learns to generalize across the teacher outputs, improving its ability to handle multi-class vulnerability detection tasks. Han et al. (Han and Zheng, 2025) follow this approach with an enhanced memory mechanism. Panichella (Panichella, 2025) combines metamorphic testing with many-objective optimization for distillation of LLMs for code, measuring model robustness across semantically equivalent code variants.
Retraining a model on all known vulnerabilities each time a new vulnerability is disclosed would demand considerable computational resources and time. Continual Learning (T2.3.5) enables models to incrementally learn new knowledge over time, such as new programming languages or up-to-date vulnerability type data (Wei et al., 2025), without forgetting what has already been learned (Catastrophic Forgetting). Gao et al. (Gao et al., 2023) propose an adaptive regularization strategy that preserves the most representative (i.e., informative and diverse) samples in each dataset for model retraining under project-level, language-level, and time-level continual learning settings. Similarly, Zhang et al. (Zhang et al., 2024c) present an incremental learning approach that allows the model to incorporate new programming languages without degrading performance on those it has already learned. Tian et al. (Tian et al., 2025) introduce a training-free parameter fusion method that merges independent classification heads fine-tuned on different vulnerability types, allowing the fused model to adapt to new categories while preserving multi-class classification capabilities.
Reinforcement Learning (RL) (T2.3.6) introduces a reward-based feedback loop, allowing models to optimize sequential decision-making processes. Several studies focus on refining the model’s reasoning capabilities. Weyssow et al. (Weyssow et al., 2025) employ Reinforcement Learning from AI Feedback (RLAIF) to distill structured reasoning, reinforcing informative and coherent while penalizing misleading reasoning. Similarly, Simoni et al. (Simoni et al., 2025) integrate an RL-based loss function to fine-tune instruction-based LLMs. Ren et al. (Ren et al., 2024) introduce a reward mechanism with prompt tuning that encourages the model to discover prompts that maximize detection performance. Beyond reasoning, RL is utilized to optimize input and feature extraction. Jiang et al. (Jiang et al., 2025a) propose an RL framework to learn vulnerability-relevant structures within code snippets. Zhao et al. (Zhao and Guan, 2025) use RL to generate high-quality adversarial examples, selecting optimal perturbations that minimize the model’s vulnerability prediction probability to create harder training samples.
@@
Other Data-Centric (T2.3.7) learning techniques have been explored in the surveyed studies to improve model performance, e.g., Positive and Unlabeled (PU) learning, active learning, and adversarial training.
@@
Forms of active learning aim to reduce labeling costs by strategically selecting the most informative samples for annotation (Alrashedy et al., 2024; Lan et al., 2025; Cheng et al., 2025). For example, Lan et al. (Lan et al., 2025) combine dataset maps with active learning to prioritize valuable samples while filtering out those that may harm performance.
Similarly, in self-paced learning, the most suitable data is selected for each training epoch, from easy to hard learning samples (Cheng et al., 2025).